OpenStudy (anonymous):

Choose the correct product of (7x − 6)2. 49x2 + 84x + 36 49x2 − 36 49x2 + 36 49x2 − 84x + 36

OpenStudy (anonymous):

i got 49x^2-36..is that right?

OpenStudy (cwrw238):

(7x − 6)2. = 7x(7x - 6) - 6(7x - 6)

OpenStudy (anonymous):

No its the last one

OpenStudy (cwrw238):

= 49x^2 - 42x - 42x + 36

OpenStudy (anonymous):

how @Ashley1997 ? could one of you show me? please?

OpenStudy (cwrw238):

can you complete this?

OpenStudy (cwrw238):

whats -42x - 42x?

OpenStudy (anonymous):

I think its the last one

OpenStudy (anonymous):

@cwrw238 -84

OpenStudy (cwrw238):

don't forget the x - its -84x

OpenStudy (anonymous):

but where did you get the 42 from...?

OpenStudy (cwrw238):

have a look at my first post

OpenStudy (anonymous):

@zakiaboo That is my best guess. because thats the answer or closest answer

OpenStudy (cwrw238):

7x * -6 and -6 * 7x

OpenStudy (anonymous):

@ceit yes i kno thats the right answer but @cwrw238 is showing me how to do it..so I can learn how to actually do it lol :)

OpenStudy (anonymous):

@cwrw238 ohhh so you have to multiply 7*6 first?

OpenStudy (anonymous):

@zakiaboo Oh sorry I am learning how to draw things out. I just started here so yeah

OpenStudy (anonymous):

Its cool :)

OpenStudy (cwrw238):

yes 7x( 7x - 6) = 7x*7x + x*-6 = 49x^2 - 42x

OpenStudy (anonymous):

hmm okay let me take notes.

OpenStudy (cwrw238):

a simple example of multiplying 2 binomials is (x + y)(x + y): first multiply (x+y) in second bracket by x then by y. = x(x + y) + y(x + y) = x^2 + xy + xy + y^2 = x^2 + 2xy + y^2

OpenStudy (anonymous):

\[49x^2-84x+36\]\[(-6+7x)^2=(-6+7x)(-6+7x)\]\[-6 (-6+7 x)+7 x (-6+7 x) \] \[-6(-6+7x)+7(-6x+7x^2)\] \[-6(-6)=36 \] \[-6(7)+7(-6)=-84\] \[7(7x^2)=49x^2\] \[formula=ax^2+bx+c\] \[49x^2-84x+36\]